Circular No. 8273

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004K IN ESO 579-G22 Further to IAUC 8272, J. Graham and W. Li report the LOSS dis-

covery, on KAIT images taken on Jan. 19.6 (mag ∼ 17.2) and 21.5 UT (mag ∼ 17.1), of an apparent supernova located at α = 14^{h}23^{m}39^{s}.85, δ = −19^{o}26′50′′.0 (equinox 2000.0), which is 15′′.5 west and 11′′.8 south of the nucleus of ESO 579-G22. A KAIT image taken on 2003 July 1.2 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2004G IN NGC 5668 N. Elias-Rosa, Osservatorio Astronomico di Padova (OAP); G. Pig-

nata, European Southern Observatory (ESO); S. Benetti, G. Blanc, A. Della Valle, A. Pastorello, G. Altavilla, H. Navasardyan, M. Turatto, and L. Zampieri, OAP; E. Cappellaro, Osservatorio Astronomico di Capodimonte; and F. Patat, ESO, on behalf of the ERTN (IAUC 7987), report, from inspection of spectra obtained on Jan. 21.15 UT with the Calar Alto 2.2-m reflector (+ CAFOS; range 380–880 nm; resolution 1.1 nm), and Jan. 21.18 with the Asiago 1.8-m telescope (+ AFOSC; range = 355–780 nm, resolution 2.4 nm), that SN 2004G (IAUC 8272) is a type-II supernova, ∼ 5 months after explosion. The expansion velocities deduced from the Hα and Hβ absorptions are ∼ 6600 and 4300 km/s, respectively.

SUPERNOVA 2004H IN IC 708 C. L. Gerardy and G. H. Marion, University of Texas at Austin, report

that a low-resolution optical spectrogram (resolution 600; range 430–700 nm) of SN 2004H (cf. IAUC 8272), obtained on Jan. 22.35 UT with the 2.7-m Harlan J. Smith Telescope (+ LCS spectrograph) at McDonald Observatory, shows it to be a type-Ia supernova near maximum light; adopting the 9497 km/s redshift of IC 708 from the NASA/IPAC Extragalactic Database, the expansion velocity of the Si II feature (rest wavelength 635.5 nm) is 10600 km/s. The 580-nm Si II absorption feature appears unusually strong relative to the 615-nm feature, indicating that SN 2004H is likely a subluminous type-Ia object.

COMET C/2001 Q4 (NEAT) Total visual magnitude estimates: 2003 Dec. 22.09 UT, 9.8 (A. Amo-

rim, Florianopolis, Brazil, 0.14-m refl.); 2004 Jan. 10.46, 9.0 (D. A. J. Seargent, The Entrance, N.S.W., 10×50 binoculars); 14.02, 9.0 (Amorim).

2004 January 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green